# Prediction Sets and Conformal Inference with

### v2

@@ Appendix B Uniform consistency of the kernel smoothing estimator
Similarly define the population version $P_{f}(x)=\operatorname{\mathrm{E}}\left[f(Y^{L},Y^{U};(a_{m},b_{m})_{m=1}^{M})\right]$. It is straightforward to show that $\mathcal{F}_{M}$ is a measurable VC class of functions, with a constant enevelop function $F=1$ and hence for any probability measure $Q$ over the support of $(Y^{L},Y^{U})$,
@@
where $C,v$ depends only on the VC characterisitcs of $\mathcal{F}_{M}$ and $r\geq 1$ vaart2013WeakConvergence. Then the next lemma follows from Corollary 1 in einmahl2000EmpiricalProcess with only minor modifications under some conditions on the smoothing kernel and bandwidth sequence.
@@
Suppose that $K(\frac{x-X}{h})=\prod_{j=1}^{d}K_{0}\left(\frac{x^{j}-X^{j}}{h}\right)$ where $x^{j}$ denote the $j$-th component of $x$, and $K_{0}$ is conitnuous over a compact support, $\int K(s)\,\mathrm{d}s=1.$
@@
$\sup_{x\in\mathcal{X}^{\prime}}\sup_{f\in\mathcal{F}_{M}}\left\lvert P_{n,f}(x)-\bar{P}_{n,f}(x)\right\rvert=O\left(\sqrt{\frac{\log\left\lvert h\right\rvert}{nh^{d}}}\right),$
@@
$\sup_{x\in\mathcal{X}^{\prime}}\sup_{f\in\mathcal{F}_{M}}\left\lvert P_{n,f}(x)-P_{f}(x)\right\rvert=O\left(\sqrt{\frac{\log\left\lvert h\right\rvert}{nh^{d}}}+h^{2}\right).$ (17)
As shown in hansen2008UniformConvergence Theorem 8, we can let $\mathcal{X}^{\prime}=\mathcal{X}_{n}$ be an expanding subsets at a suitable rate, and the uniform convergence rate needs to be multiplied by a factor of $\delta_{n}^{-1}$ where $\delta_{n}=\inf_{x\in\mathcal{X}_{n}}p_{X}(x)$.
@@ Appendix C Differential conformal procedure
    $\tilde{C}_{I}(x)=\left[\hat{\tau}_{0}(x)-\vartheta,\hat{\tau}_{1}(x)+\vartheta\right],\quad x\in\mathcal{X}.$ (18)
It can be seen that in the second step, the endpoints $\hat{\tau}_{0}(x)$ and $\hat{\tau}_{1}(x)$ are treated symmetrically. We could instead consider a differentiated approach on the two endpoints of $\hat{C}_{I}$. Define the following contour set $\mathcal{W}=\left\{(w_{0},w_{1}):P_{\mathcal{I}_{2}}(s_{j}(\hat{\tau}_{0}-w_{0},\hat{\tau}_{1}+w_{1}))\gtrsim 1-\alpha\right\}$. And we can pick $(w_{0}^{*},w_{1}^{*})=\min_{\mathcal{W}}\left\lVert(w_{0}^{*},w_{1}^{*})\right\rVert$.
@@
$\min\left\lVert w\right\rVert\quad\text{s.t.}\quad P_{\mathcal{I}_{2}}\left(s(Y_{j}^{L},Y_{j}^{U};\tilde{C}_{M})\leq 0\right)\gtrsim 1-\alpha.$ (19)
@@ Appendix D Metrics for sets
$d_{H}(A,B)=\max\left\{\sup_{a\in A}\inf_{b\in B}\left\lvert a-b\right\rvert,\sup_{b\in B}\inf_{a\in A}\left\lvert a-b\right\rvert\right\}.$ (20)
@@
However, when the prediction set is a union of intervals, these two metrics are not equivalent. In fact, consider $A\subset\mathbb{R}$ and $B=A\cup\left\{b\right\}$, for some $b\in A^{c}$ such that $d(b,A)>0$, then the Hausdorff distance between $A$ and $B$ is $d_{H}(A,B)=d(b,A)>0$, while the volume of the symmetric difference is 00. Thus, the volume of symmetric difference has the advantage of being less sensitive to a difference of negligible sets and seems more approporiate for practical prediction sets estimations.
@@ Appendix E Estimation and conformal inference based on quantile regression
$P\left(Y^{L}\geq q^{L}_{\frac{\alpha}{2}}(x),Y^{U}\leq q^{U}_{1-\frac{\alpha}{2}}(x)\mid X=x\right)\geq 1-\alpha$ (21)
and $\hat{C}_{q}(x)=[\hat{q}^{L}_{\frac{\alpha}{2}}(x),\hat{q}^{U}_{1-\frac{\alpha}{2}}(x)]$ as an estimator for $C_{q}(x)=[{q^{L}_{\frac{\alpha}{2}(x)},q^{U}_{1-\frac{\alpha}{2}}(x)}]$ is an asymptotically valid, albeit not necessarily optimal, prediction set. Here $\hat{q}^{L}_{\frac{\alpha}{2}}(x),\hat{q}^{U}_{1-\frac{\alpha}{2}}(x)$ are the consistent quantile regression estimators. There is a vast body of research on quantile regression, see koenker2017QuantileRegression, athey2019GeneralizedRandom for random forest based quantile regression and li2021NonparametricQuantile for nonparametric quantile regression with mixed continuous and discrete data.
@@
$\tilde{C}_{q}(x)=\left[\hat{q}^{L}_{\frac{\alpha}{2}}(x)-\vartheta^{q}_{1-\alpha},\hat{q}^{U}_{1-\frac{\alpha}{2}}(x)+\vartheta^{q}_{1-\alpha}\right].$ (22)

### v3

@@ Appendix B Uniform consistency of the kernel smoothing estimator
Similarly define the population version $P_{f}(x)=\operatorname{\mathrm{E}}\left[f(Y^{L},Y^{U};(a_{m},b_{m})_{m=1}^{M})\right]$. It is straightforward to show that $\mathcal{F}_{M}$ is a measurable VC class of functions, with a constant envelop function $F=1$ and hence for any probability measure $Q$ over the support of $(Y^{L},Y^{U})$,
@@
where $C,v$ depends only on the VC characteristics of $\mathcal{F}_{M}$ and $r\geq 1$ (van der Vaart and Wellner,, 2013). Then the next lemma follows from Corollary 1 in Einmahl and Mason, (2000) with only minor modifications under some conditions on the smoothing kernel and bandwidth sequence.
@@
Suppose that $K(\frac{x-X}{h})=\prod_{j=1}^{d}K_{0}\left(\frac{x^{j}-X^{j}}{h}\right)$ where $x^{j}$ denote the $j$-th component of $x$, and $K_{0}$ is continuous over a compact support, $\int K(s)\,\mathrm{d}s=1.$
@@
$\sup_{x\in\mathcal{X}^{\prime}}\sup_{f\in\mathcal{F}_{M}}\left\lvert P_{n,f}(x)-\bar{P}_{n,f}(x)\right\rvert=O\left(\sqrt{\frac{\left\lvert\log h\right\rvert}{nh^{d}}}\right),$
@@
$\sup_{x\in\mathcal{X}^{\prime}}\sup_{f\in\mathcal{F}_{M}}\left\lvert P_{n,f}(x)-P_{f}(x)\right\rvert=O\left(\sqrt{\frac{\left\lvert\log h\right\rvert}{nh^{d}}}+h^{2}\right).$ (19)
As shown in Hansen, (2008) Theorem 8, we can let $\mathcal{X}^{\prime}=\mathcal{X}_{n}$ be an expanding subsets at a suitable rate, and the uniform convergence rate needs to be multiplied by a factor of $\delta_{n}^{-1}$ where $\delta_{n}=\inf_{x\in\mathcal{X}_{n}}p_{X}(x)$.
@@ Appendix C Differential conformal procedure
    $\tilde{C}_{I}(x)=\left[\hat{\tau}_{0}(x)-\vartheta,\hat{\tau}_{1}(x)+\vartheta\right],\quad x\in\mathcal{X}.$
It can be seen that in the second step, the endpoints $\hat{\tau}_{0}(x)$ and $\hat{\tau}_{1}(x)$ are treated symmetrically. We could instead consider a differentiated approach on the two endpoints of $\hat{C}_{I}$. Define the following contour set
$\mathcal{W}=\left\{(w_{0},w_{1}):P_{\mathcal{I}_{2}}(s_{j}(\hat{\tau}_{0}-w_{0},\hat{\tau}_{1}+w_{1}))\geq(1-\alpha)(1+\left\lvert\mathcal{I}_{2}\right\rvert^{-1})\right\}.$
And we can pick $(w_{0}^{*},w_{1}^{*})=\arg\min_{\mathcal{W}}\left\lVert(w_{0}^{*},w_{1}^{*})\right\rVert$.
@@
$\min\left\lVert w\right\rVert\quad\text{s.t.}\quad P_{\mathcal{I}_{2}}\left(s(Y_{j}^{L},Y_{j}^{U};\tilde{C}_{M})\leq 0\right)\geq(1-\alpha)(1+\left\lvert\mathcal{I}_{2}\right\rvert^{-1}).$
@@ Appendix D Metrics for sets
$d_{H}(A,B)=\max\left\{\sup_{a\in A}\inf_{b\in B}\left\lvert a-b\right\rvert,\sup_{b\in B}\inf_{a\in A}\left\lvert a-b\right\rvert\right\}.$
@@
However, when the prediction set is a union of intervals, these two metrics are not equivalent. In fact, consider $A\subset\mathbb{R}$ and $B=A\cup\left\{b\right\}$, for some $b\in A^{c}$ such that $d(b,A)>0$, then the Hausdorff distance between $A$ and $B$ is $d_{H}(A,B)=d(b,A)>0$, while the volume of the symmetric difference is 00. Thus, the volume of symmetric difference has the advantage of being less sensitive to a difference of negligible sets and seems more appropriate for practical prediction sets estimations.
@@ Appendix E Estimation and conformal inference based on quantile regression
$P\left(Y^{L}\geq q^{L}_{\frac{\alpha}{2}}(x),Y^{U}\leq q^{U}_{1-\frac{\alpha}{2}}(x)\mid X=x\right)\geq 1-\alpha$
and $\hat{C}_{q}(x)=[\hat{q}^{L}_{\frac{\alpha}{2}}(x),\hat{q}^{U}_{1-\frac{\alpha}{2}}(x)]$ as an estimator for $C_{q}(x)=[{q^{L}_{\frac{\alpha}{2}(x)},q^{U}_{1-\frac{\alpha}{2}}(x)}]$ is an asymptotically valid, albeit not necessarily optimal, prediction set. Here $\hat{q}^{L}_{\frac{\alpha}{2}}(x),\hat{q}^{U}_{1-\frac{\alpha}{2}}(x)$ are the consistent quantile regression estimators. There is a vast body of research on quantile regression, see Koenker, (2017), Athey et al., (2019) for random forest based quantile regression and Li et al., (2021) for nonparametric quantile regression with mixed continuous and discrete data.
@@
$\tilde{C}_{q}(x)=\left[\hat{q}^{L}_{\frac{\alpha}{2}}(x)-\vartheta^{q}_{1-\alpha},\hat{q}^{U}_{1-\frac{\alpha}{2}}(x)+\vartheta^{q}_{1-\alpha}\right].$